# Graph Neural Networks for Binary Programming

### v1

@@ 4.2 Properties of the BP Problem
Consider the Ising model with constant coefficients (Dziarmaga, 2005; Duminil-Copin & Tassion, 2016), that reads:
@@
To further analyze the solution of the problem in a more general manner, we start by proving the following theorem:
Theorem 4.2.
The BP problem in Equation 1, at its optimum, i.e., $f({\bf x}_{o})$, is piece-wise differentiable with respect to changes of the observed vector ${\bf b}$.
@@
where $\|{\boldsymbol{\delta}}\|=1$ is a direction vector. Let ${\bf x}_{\epsilon}$ be the minimizer of Equation 4 and let $f_{\epsilon}=f({\bf x}_{\epsilon},\epsilon)$. Then, by the definition of the minimizer, we have that for any $\epsilon$
@@
$f_{\epsilon}$ $=$ $f({\bf x}_{\epsilon},\epsilon)$
$\leq$ $f({\bf x}_{0},\epsilon)=f({\bf x}_{0},0)+\epsilon{\boldsymbol{\delta}}^{\top}{\bf x}_{0}=f_{0}+\epsilon{\boldsymbol{\delta}}^{\top}{\bf x}_{0}.$
@@
Because the norms of ${\bf x}_{0}$ and ${\bf x}_{\epsilon}$ are bounded (as they are binary) we obtain upper and lower bounds for the change in the optimal solution in the direction ${\boldsymbol{\delta}}$. This is the definition of Clarke derivative (Michel & Penot, 1992), which extends the definition of the derivative beyond a point to an interval. The function $f({\bf b})$ is therefore piece-wise linear. ∎
Theorem 4.2 shows that the BP problem in Equation 1 is piece-wise differentiable with respect to ${\bf b}$, a property that we now utilize in order to characterize the BP solution behavior with respect changes in the observed vector ${\bf b}$.
Observation 4.3.
The BP solution is piece-wise constant with respect to changes in ${\bf b}$.
@@
To study the effect of a change in ${\bf b}$ on the solution ${\bf x}_{\epsilon}$ we now choose ${\boldsymbol{\delta}}={\bf e}_{j}$, that is, the unit vector in direction $j$. This simplifies Equation Equation 8 to
@@
Furthermore, in this case $f_{0}=f_{\epsilon}$ which implies that ${\bf x}_{0}={\bf x}_{\epsilon}$ is a solution of the perturb problem. Thus we obtain that
@@
Similarly, if ${\bf x}_{0_{j}}=1$ and $\epsilon<0$ and ${\bf x}_{\epsilon_{j}}\not={\bf x}_{\epsilon_{j}}$, then ${\bf x}_{\epsilon_{j}}=0$. But this again yields a contradiction which means
@@
This implies that the function $q(\cdot)$ is piece-wise constant with respect to changes in ${\bf b}$. ∎
Below, in Example 2, we provide a concrete example of this property, and showcase its piece-wise constant landscape behavior with respect to the vector ${\bf b}$.
@@
In Figure 1 we plot $\phi$ as a function of $s$ and $t$ for four different experiments where the matrix, ${\bf A}$, the vector, ${\bf b}$ and the perturbations ${\bf b}_{1}$ and ${\bf b}_{2}$ are randomly chosen. In all cases, we see that the solution is constant over a range of values and then sharply changes. Such surfaces are good candidates for deep network approximation.
@@
Figure 1: The function $\phi(s,t)$ in Equation 13 for four different random realizations of ${\bf A},{\bf b},{\bf b}_{1}$ and ${\bf b}_{2}$. $\phi(s,t)$ is piece-wise constant, favored by neural networks.
5 Approximating BP Solutions with GNNs
5.1 The Heterophilic Behavior of BP Solutions
While it is difficult to predict the exact distribution of the solution ${\bf x}$, a few simple cases have been studied in the context of the Ising model (Dziarmaga, 2005). For this model, it is known that when the solution resides on a regular lattice, it can exhibit high-frequency patterns. In graphs and GNNs, such solutions are commonly affiliated with heterophily (Pei et al., 2020; Choi et al., 2023). To demonstrate that, we solve four different problems on a lattice of size $32\times 32$ using the SAB algorithm (Goto et al., 2019) and plot the solutions in Figure 2.
@@ 5.1 The Heterophilic Behavior of BP Solutions
It is observed that the obtained solutions are highly irregular and contain small connected components of uniform values. This implies that typical graph diffusion, or smoothing-based GNNs may face difficulties in expressing such non-smooth patterns. Therefore, as we show in Section 5.2, we leverage the power of heterophily-designated GNNs, combined with a specialized BP-aware component.
5.2 Binary Programming Graph Neural Network
In Section 5.1 and Figure 2 we established that the spatial behavior of the solution of the BP problem, ${\bf x}_{0}$, is heterophilic. In this section, we design a GNN that is capable of generating spatially non-smooth patterns while being aware of its application – predicting approximations of BP solutions.
GNN Backbone. In recent years, many heterophily-designated GNN architectures have been proposed. A non-exhaustive list includes (Zhu et al., 2020; Luan et al., 2022; Rusch et al., 2022; Eliasof et al., 2023b; Wang et al., 2022; Choi et al., 2023; Zhao et al., 2023; Rossi et al., 2023). In this paper, we chose to use a reaction-diffusion based network, which is similar to ACMP (Wang et al., 2022) and GREAD (Choi et al., 2023). Our motivation for following this approach is that it is inspired by Turing patterns (Turing, 1952), which are complex and have relation to the Ising model (Merle et al., 2019), which is an instance of a BP problem, as discussed in Example 1. We note that while other heterophilic designated architectures can be used, this is not the focus of our work. Rather, our contribution is drawing the link between GNNs, node classification, and the solution of Binary Programming problems.
QUBO Objective Awareness. While the chosen reaction-diffusion GNN is powerful (also as shown in our experiment in Section 7.2), we argue that it is important also to allow the network to be aware of the objective we seek to minimize, that is, the QUBO function, as shown in Equation 1. To this end, note that the QUBO objective can be written as the sum of the nodal residual of the vector ${\bf r}={\bf x}\odot({\bf A}{\bf x}+{\bf b})$, where $\odot$ is the Hadamard product. Thus, we add the following term to the node features at each layer:
@@ 5.2 Binary Programming Graph Neural Network
This modification allows the network to implicitly track the nodal QUBO loss, allowing the network to be aware of the optimization problem at hand. As we show in Section 7.2, this addition improves the solution approximation and training convergence, as we show in our experiments.
BPGNN. Given input features ${\bf b}$, we first embed them to obtain the initial features using a Multilayer Perceptron (MLP) encoder denoted by ${\rm{enc}}$, as follows:
@@
As described in Equation 14, we also include QUBO objective awareness. Thus, a BPGNN layer reads:
@@
where $g$ is an MLP. As discussed earlier, for our $\rm{GNN}$ implementation we use a reaction-diffusion GNN, discussed In Appendix A. Since we consider a binary problem, after $L$ layers, we use a decoder (classifier) denoted ${\rm{dec}}$, implemented by a linear layer to obtain the predicted solution:
@@ 6 Generating Training Data
In order to train the network, given a matrix ${\bf A}$, we require to have observation-solution pairs $\{{\bf b}^{(j)},{\bf x}_{{o}}^{(j)}\}$, $j=0,\ldots,n-1$. Given these pairs, we use the binary cross-entropy loss function between ${\bf x}_{p}^{(j)}$, the predicted solution obtained from BPGNN, and the ground-truth solution ${\bf x}_{o}^{(j)}$.
Data Acquisition is Computationally Expensive. An important aspect is how data pairs observation-solution pairs $\{{\bf b}^{(j)},{\bf x}_{{o}}^{(j)}\}$ can be obtained. Because, as discussed throughout the paper, given an observation vector ${\bf b}^{(j)}$, finding the optimal solution ${\bf x}_{{o}}^{(j)}$ is an NP-complete problem. For small problems (e.g., $k\leq 20$) one can use exhaustive search techniques. This approach requires solving the optimization problem in order to obtain a training set. Clearly, such an approach cannot be used for large-scale problems. For example, if $k=1024$, one has to check $2^{1024}$ possible solutions.
Practical Data Pairs Generation. In light of the computational challenge described above, we derive a novel data generation technique. Specifically, instead of choosing the observed vector ${\bf b}^{(j)}$ and then solving an NP-complete problem to find ${\bf x}_{{o}}^{(j)}$, we choose first ${\bf x}_{{o}}^{(j)}$, and only then find a corresponding ${\bf b}^{(j)}$. This procedure allows to create data pairs while avoiding the solution of an NP-complete problem. Specifically, we consider a relaxed optimization problem, where ${\bf x}$ is continuous in the interval $[0,1]$:
@@
One popular way to solve this minimization problem is by the log barrier approach (Nocedal & Wright, 1999), which is obtained by adding the logarithmic terms as follows:
@@
As shown in Nocedal & Wright (1999), the conditions for a minimum are:
@@
${\bf b}_{o}$ $=$ $-({\bf A}+{\bf A}^{\top}){\bf x}_{o}+{\frac{\mu}{{\bf x}_{o}}}-{\frac{\mu}{1-{\bf x}_{o}}},$ (22)
${\bf b}$ $=$ ${\bf b}_{o}+\sigma^{2}{\bf z},$ (23)
where ${\bf z}\sim N(0,{\bf I})$ is a random vector. The vector ${\bf b}_{o}$ corresponds to an approximate solution to the optimization problem in Equation 19, which is also the solution of the binary problem in Equation 1. By adding a noise term $\sigma^{2}{\bf z}$ to the observed vector in Equation 23, the solution of the continuous optimization problem is now different from the solution of the discrete binary problem. Nonetheless, as explored in the previous section, the solution of the discrete problem varies little when $\sigma$ is not large. We therefore use a small number of Tabu search steps, starting from ${\bf x}_{o}$ to modify the solution if needed. While there is no guarantee that the solution is minimal, we have verified by comparison with exhaustive search solutions on small-scale problems that indeed we obtain the global minima every time. This approach, of choosing ${\bf x}_{o}$ rather than ${\bf b}$ allows us to quickly generate large amounts of data even for large-scale problems (with hundreds or thousands of decision variables), where finding the exact solution is computationally intractable. Our procedure for data generation is summarized in Algorithm 1.
@@
4:  Use up to ten Tabu search iterations to solve Equation 1 starting from ${\bf x}_{o}$ to obtain improved an solution ${\bf x}$.
@@ 7 Numerical Experiments
Method $\downarrow$ / Problem Size $\rightarrow$ Small ($k=10$) Medium ($k=256$) Large $(k=1024)$
Acc(%) Rel. QUBO Inference Acc(%) Rel. QUBO Inference Acc(%) Rel. QUBO Inference
Objective Time (ms) Objective Time (ms) Objective Time (ms)
Naive approach.
Exhaustive Search 100 0 8 100 0 $10^{7}$ days 100 0 $10^{302}$ days
Classical solvers.
@@
Table 1: A quantitative comparison between a naive exhaustive search method, classical methods, GNNs, and our BPGNN.
@@
In this section, we evaluate the performance of BPGNN and other baselines to understand whether our theoretical observations regarding the behavior of the BP, characterized in Section 4.2 and Section 5.1, hold in practice.
@@
Baselines. We consider several baselines in our experiments: (i) naive methods, i.e., exhaustive search, (ii) classical solvers such as TS (Glover & Laguna, 1998) and SAB (Goto et al., 2019), and (iii) GNN baselines, namely GCN (Kipf & Welling, 2016), GAT (Veličković et al., 2018), GCNII (Chen et al., 2020), and Dir-GNN (Rossi et al., 2023). We chose to compare our method with those baselines as they are popular and represent methods from recent years in the context of GNNs. Also, Dir-GNN is heterophily designated, thus suitable to be tested with the heterophilic behavior of BP solutions, as discussed in Section 5.1.
Metrics. In our experiments, we consider three important metrics: (i) the solution accuracy (higher is better), (ii) the relative QUBO objective value (lower is better), and (iii) the computation time required to achieve the approximated solution. We elaborate on the metrics in Appendix B.
Datasets. Our experiments include three datasets generated by our procedure described in Section 6, in varying sizes - small ($k=10$), medium ($k=256$) and large ($k=1024$). Note that the settings of $k=\{256,1024\}$ are considered highly challenging, as there are $2^{\{256,1014\}}$ possible solutions to the binary programming problem. We provide further details on the datasets in Appendix C.
Implementation. Our code is implemented in Pytorch (Paszke et al., 2019) and Pytorch-Geometric (Fey & Lenssen, 2019).
Inference. We consider two main inference modes to obtain an approximation of the solutions of BP problems.
  • Neural. As standard in neural networks, we employ BPGNN to approximate the QUBO solution.
  • Hybrid. We obtain a solution approximation using BPGNN. This solution is considered an initial point in a classical solver. In our experiments, we used a small fixed number (up to 10) iterations of TS (Glover & Laguna, 1998), as discussed in Section 3.1.
7.1 QUBO Solution Approximation
We now compare the obtained solution using the various baselines, as previously discussed. Our results are reported in Table 1. We report results on small, medium, and large datasets. As discussed in 2, here we consider the case where the adjacency matrix ${\bf A}$ from Equation 1 is constant, while the observed vector ${\bf b}$ is variable. To ensure the statistical significance of our results, for each problem size, we generate 10 instances of the problem matrix ${\bf A}$ and report the average and standard deviation metrics in Table 1 and E.1.
Our results indicate that: (i) Naturally, an exhaustive search yields an accurate solution. However, it is not practical for medium or large problems (i.e., as the number of decision variables $k$, or nodes in the context of GNNs, grows). (ii) Classical solvers achieve impressive accuracy on small problems. However, the accuracy degrades as the problem grows, and the required computational time significantly grows. (iii) GNNs are capable of modeling the considered QUBO problem, effectively learning the discussed map $q(\cdot)$ in Section 4. We also see that in accordance with our understanding regarding the behavior of BP solutions, heterophily designated GNNs such as Dir-GNN and our BPGNN offer higher accuracy and a lower relative QUBO objective compared to other GNNs. (iv) Combining neural approaches, such as our BPGNN, together with classical solvers as TS, offers improved performance, at a fraction of the cost of approximating the problem solution with classical solvers from scratch. Therefore, we deem that our approach can also serve as a strong prior to initializing classical solvers.
7.2 QUBO Aware Features Improve Performance
As presented in Equation 14, our BPGNN includes QUBO-aware node features, which is a novel component of our approach. Therefore, we now examine its impact on the performance. In Table 2, we report the results. It can be seen that including QUBO-aware features adds negligible computation time (less than a millisecond) while improving solution accuracy, on average, by $1.7\%$ and reducing the relative QUBO objective by a factor of almost 2, and improving training convergence, as we show in Section E.2.
QUBO Features Large $(k=1024)$
Acc(%) Rel QUBO Inference
Objective Time (ms)
[…]
Table 2: The impact of adding the QUBO aware features from Equation 14 to the node features.
8 Limitations and Open Challenges
The experiments above demonstrate our ability to train a network to solve the QUBO problem Equation 1 for a fixed matrix ${\bf A}$ and a variable vector ${\bf b}$. Our theoretical observations and empirical findings pave the way to further research in the connection between solving complex combinatorial problems such as BP problems using GNNs. We now provide several possible significant extensions of our work, both theoretically and in practical applications.
Variable Adjacency Matrix ${\bf A}$. A first open challenge is to extend this work to the case where the matrix ${\bf A}$ is also allowed to change, which is an important problem, see for example (Oshiyama & Ohzeki, 2022) for non-neural solutions. This aspect is important because it allows to model different underlying problems and graphs, rather than only a variable observed vector ${\bf b}$.
Constrained Problems. In this paper, we consider QUBO, which is a subset of unconstrained binary optimization problems. However, many binary problems require considering constraints. Several examples are the knapsack problem (Kellerer et al., 2004), topology optimization (Sivapuram & Picelli, 2018), and protein design (Strokach et al., 2020).
Nonquadratic Problems. In this paper, we study the approximation of QUBO problems using GNNs, but future research can also consider nonquadratic problems, for example, (Murray & Ng, 2010) used in applications such as protein design (Allouche et al., 2014) and telecommunication network optimization (Resende & Pardalos, 2008).
9 Summary
This paper introduces a novel approach, merging graph learning and optimization problem-solving, with an innovative perspective on approximating combinatorial problems solutions by using GNNs. We analyze the input-output relationship of QUBO problems and show that their sensitivity is piecewise constant, and their solutions are heterophilic. These insights guide the design of BPGNN, a GNN architecture to approximate input-solution mappings.
To efficiently generate input-solution data pairs, we propose a novel and scalable mechanism that enables data generation while avoiding the solution of the NP-complete QUBO problem from scratch. Our Numerical results affirm the superiority of BPGNN in accuracy for fixed graph problems, outperforming other GNN solutions while improving computational efficiency compared to classical solvers.
Furthermore, this work not only establishes and evaluates the link between GNNs and BP problems, but also outlines promising future research directions in this emerging research area.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. We believe it draws an important link between the approximation of the solution of NP-complete problems and neural networks, specifically by utilizing Graph Neural Networks. This work is a first step in this direction and we discuss important future directions. The success and growth of this research front can positively impact various fields, from bioinformatics, to finance, logistics, and more. We are not aware of negative societal consequences of our work.
Appendix A BPGNN Architecture
We now provide details on the implementation of our BPGNN, which combines concepts of reaction-diffusion similar to (Wang et al., 2022; Choi et al., 2023) and our proposed QUBO aware node features.
Overall, our BPGNN is defined by the following ordinary differential equations (ODE) and initial conditions:
${\bf r}({\bf h},{\bf A},{\bf b};t)=q({\bf h}\odot({\bf A}{\bf h}+{\bf b});t)$ (25)
${\bf h}(t=0)={\bf h}_{0}={\rm{enc}}({\bf b})$ (26)
Here, ${\bf h}(t)\in\mathbb{R}^{k\times d}$ is a hidden feature matrix with $d$ features that depends on the initial state ${\bf h}_{0}$, and the input node features ${\bf b}\in\mathbb{R}^{k\times 1}$. ${\dot{\bf h}=\frac{d{\bf h}(t)}{dt}}$ denotes the first-order derivative with respect to time. ${\bf L}$ is the graph Laplacian as in (Eliasof et al., 2021). ${\bf r}$ is a continuous version of the QUBO objective aware features as shown in Equation 14, and $q$ is an MLP. The initial state is obtained by embedding the input node features using a single layer MLP; ${\bf h}_{0}={\rm{enc}}({\bf b})$. Also, ${\bf\Sigma}(t)\geq 0$ is a diagonal matrix with non-negative learnable diffusion coefficients on its diagonal. The reaction term $f$ from Equation (24) is a learnable point-wise non-linear function that depends on the current hidden state ${\bf h}(t)$.
In order to obtain a graph neural layer from Equation (24)-(26), we use the forward Euler discretization method, as is common in GNNs (Eliasof et al., 2021; Chamberlain et al., 2021; Gravina et al., 2022). Therefore, a BPGNN layer is defined by the following update rules shown in Algorithm 2.
Algorithm 2 BPGNN Layer
Input: Node features ${\bf h}^{(\ell)}\in\mathbb{R}^{k\times d}$
Output: Updated node features ${\bf h}^{(l+1)}\in\mathbb{R}^{k\times d}$
4:  Return ${\bf h}^{(l+1)}$.
Appendix B Metrics
We now discuss the three metrics reported in our experiments, in Section 7.

### v2

@@ 4.2 Properties of the QUBO Problem
Consider the Ising model with constant coefficients [10, 11], that reads:
@@
To analyze the solution in a more general manner, we start by proving the following theorem:
Theorem 2.
The QUBO problem in Equation˜1, at its optimum, i.e., $f({\bf x}_{o})$, is piece-wise differentiable with respect to changes of the observed vector $\bf b$}.
@@
where $\|{\boldsymbol{\delta}}\|=1$ is a direction vector. Let ${\bf x}_{\epsilon}$ be the minimizer of Equation˜4 and let $f_{\epsilon}=f({\bf x}_{\epsilon},\epsilon)$. Then, by the definition of the minimizer, we have that for any $\epsilon$, the following holds:
@@
$f_{\epsilon}=f({\bf x}_{\epsilon},\epsilon)\leq f({\bf x}_{0},\epsilon)=f({\bf x}_{0},0)+\epsilon{\boldsymbol{\delta}}^{\top}{\bf x}_{0}=f_{0}+\epsilon{\boldsymbol{\delta}}^{\top}{\bf x}_{0}.$ (6)
@@
Because the norms of ${\bf x}_{0}$ and ${\bf x}_{\epsilon}$ are bounded (as they are binary) we obtain upper and lower bounds for the change in the optimal solution in the direction $\boldsymbol{\delta}$}. This is the definition of Clarke derivative [33], which extends the definition of the derivative beyond a point to an interval. The function $f({\bf x}_{o})$ is therefore piece-wise linear. ∎
Theorem 2 shows that the QUBO problem in Equation˜1 is piece-wise differentiable with respect to $\bf b$}, a property that we now utilize in order to characterize the QUBO solution behavior with respect changes in the observed vector $\bf b$}.
Observation 1.
The QUBO $\bf x$} solution is piece-wise constant with respect to changes in $\bf b$}.
@@
To study the effect of a change in $\bf b$} on the solution ${\bf x}_{\epsilon}$ we now choose ${\boldsymbol{\delta}}={\bf e}_{j}$, that is, the unit vector in direction $j$. This simplifies Equation˜8 to
@@
Furthermore, in this case $f_{0}=f_{\epsilon}$ which implies that ${\bf x}_{0}={\bf x}_{\epsilon}$ is a solution of the perturbed problem. Thus, we obtain that
@@
Similarly, if ${\bf x}_{0_{j}}=1$ and $\epsilon<0$ and ${\bf x}_{\epsilon_{j}}\not={\bf x}_{\epsilon_{j}}$, then ${\bf x}_{\epsilon_{j}}=0$. But this again yields a contradiction, which means
@@
Because Equations (10)–(11) show that the change in $\bf x$} is zero when changing b in $\epsilon^{+}$ and $\epsilon^{-}$. This implies that $\bf x$} remains constant for a wide range of perturbations in $\bf b$}. Thus, by definition, $\bf x$} is piecewise constant with respect to $\bf b$}. ∎
Below, in Example˜2, we provide a concrete example of this property, and showcase its piece-wise constant landscape behavior with respect to the vector $\bf b$}.
@@
In Figure 2, we plot $\phi$ as a function of $s$ and $t$ for four different experiments where the matrix, $\bf A$}, the vector, $\bf b$} and the perturbations ${\bf b}_{1}$ and ${\bf b}_{2}$ are randomly chosen. In all cases, we see that the solution is constant over a range of values and then sharply changes. Such surfaces are good candidates for deep network approximation.
@@
Figure 1: The function $\phi(s,t)$ in Equation˜13 for four different random realizations of ${\bf A},{\bf b},{\bf b}_{1}$ and ${\bf b}_{2}$. $\phi(s,t)$ is piece-wise constant, favored by neural networks.
@@
5 Approximating QUBO Solutions with GNNs
5.1 The Heterophilic Behavior of QUBO Solutions
While it is difficult to predict the exact distribution of the solution $\bf x$}, a few simple cases have been studied in the context of the Ising model [11]. For this model, it is known that when the solution resides on a regular lattice, aside from trivial cases discussed in Appendix C.1, it can exhibit high-frequency patterns. In graphs and GNNs, such solutions are commonly affiliated with heterophily [40, 8]. To demonstrate that, we solve four different problems on a lattice of size $32\times 32$ using the SAB algorithm [20] and plot the solutions in Figure 2.
It is observed that the obtained solutions are irregular and contain small connected components of uniform values. This implies that typical smoothing-based GNNs may face difficulties in expressing such non-smooth patterns. Therefore, as we discuss in Section 5.2, we leverage the power of heterophily-designated GNNs, combined with a specialized QUBO-aware component.
5.2 QUBO Graph Neural Network
In Section 5.1 and Figure 2, we established that the spatial behavior of the solution of the QUBO problem, ${\bf x}_{0}$, is heterophilic. In this section, we design a GNN that is capable of generating spatially non-smooth patterns while being aware of its application – predicting approximations of QUBO solutions.
GNN Backbone.
Recently, several heterophily-designated GNNs have been proposed. A non-exhaustive list includes [8, 13, 14, 44]. In this paper, we choose to use a reaction-diffusion GNN, which is similar to GREAD [8], discussed in Section 5.3. Our choice is inspired by Turing patterns [52], which are complex and have relation to the Ising model [31], which is an instance of a QUBO problem, as discussed in Example 1. Note that, while other heterophilic designated architectures can be used, this is not the focus of our work. Rather, our contribution is drawing the link between GNNs, node classification, and QUBO.
QUBO Objective Awareness.
While the chosen reaction-diffusion GNN is powerful, we show that it is important also to allow the network to be aware of the objective we seek to minimize, that is, the QUBO function, as shown in Equation˜1. To this end, note that the QUBO objective can be written as the sum of the nodal residual of the vector ${\bf r}={\bf x}\odot({\bf A}{\bf x}+{\bf b})$, where $\odot$ is the Hadamard product, which is added to the features as follows:
@@ 5.2 QUBO Graph Neural Network
This modification allows the network to implicitly track the nodal QUBO loss, allowing the network to be aware of the optimization problem at hand. As we show in our experiments in Section 7, this addition improves the solution approximation and training convergence.
QUBO-GNN.
Given input features $\bf b$}, we first embed them to obtain the initial features using a Multilayer Perceptron (MLP) encoder denoted by $\rm{enc}$}, as follows:
@@
As described in Equation˜14, we include QUBO-aware features. Thus, a QUBO-GNN layer reads:
@@
where $g$ is an MLP. As discussed earlier, for our $\rm{GNN}$ implementation, we use a reaction-diffusion GNN, discussed In Section 5.3. Since we consider a binary problem, after $L$ layers, we use a decoder (classifier) denoted $\rm{dec}$}, implemented by a linear layer to obtain the predicted solution:
@@
5.3 QUBO-GNN Architecture
We now provide details on the implementation of our QUBO-GNN, which combines concepts of reaction-diffusion similar to [8] and our proposed QUBO aware node features.
Overall, our QUBO-GNN is defined by the following ordinary differential equations (ODE) and initial conditions:
${\bf r}({\bf h},{\bf A},{\bf b};t)=q({\bf h}\odot({\bf A}{\bf h}+{\bf b});t)$
${\bf h}(t=0)={\bf h}_{0}={\rm{enc}}({\bf b})$
Here, ${\bf h}(t)\in\mathbb{R}^{k\times d}$ is a hidden feature matrix with $d$ features that depends on the initial state ${\bf h}_{0}$, and the input node features ${\bf b}\in\mathbb{R}^{k\times 1}$. $\dot{\bf h}=\frac{d{\bf h}(t)}{dt}$} denotes the first-order derivative with respect to time. $\bf L$} is the graph Laplacian as in [12]. $\bf r$} is a continuous version of the QUBO objective aware features as shown in Equation˜14, and $q$ is an MLP. The initial state is obtained by embedding the input node features using a single layer MLP; ${\bf h}_{0}={\rm{enc}}({\bf b})$. Also, ${\bf\Sigma}(t)\geq 0$ is a diagonal matrix with non-negative learnable diffusion coefficients on its diagonal. The reaction term $f$ from Equation (21) is a learnable point-wise non-linear function that depends on the current hidden state ${\bf h}(t)$.
In order to obtain a graph neural layer from Equation (21)-(21), we use the forward Euler discretization method, as is common in GNNs [5, 12]. Therefore, a QUBO-GNN layer is defined by the following update rules shown in Algorithm 1.
Implementation.
Our code is implemented in Pytorch [39] and Pytorch-Geometric [16], and will be published together with the datasets upon acceptance. All the experiments were conducted on an Nvidia RTX-4090 GPU with 24GB of memory.
Algorithm 1 QUBO-GNN Layer
Input: Node features ${\bf h}^{(\ell)}\in\mathbb{R}^{k\times d}$
Output: Updated node features ${\bf h}^{(l+1)}\in\mathbb{R}^{k\times d}$
4:  Return ${\bf h}^{(l+1)}$.
@@ 6 Generating Training Data
In order to train the network, given a matrix $\bf A$}, we require to have observation-solution pairs $\{{\bf b}^{(j)},{\bf x}_{{o}}^{(j)}\}$, $j=0,\ldots,n-1$. Given these pairs, we use the binary cross-entropy loss function between ${\bf x}_{p}^{(j)}$, the predicted solution obtained from QUBO-GNN, and the ground-truth solution ${\bf x}_{o}^{(j)}$.
Data Acquisition is Computationally Expensive.
An important aspect is how data pairs observation-solution pairs $\{{\bf b}^{(j)},{\bf x}_{{o}}^{(j)}\}$ can be obtained. Because, as discussed throughout the paper, given an observation vector ${\bf b}^{(j)}$, finding the optimal solution ${\bf x}_{{o}}^{(j)}$ is an NP-complete problem. For small problems (e.g., $k\leq 20$), one can use exhaustive search techniques. This approach requires solving the optimization problem in order to obtain a training set. Clearly, such an approach cannot be used for large-scale problems. For example, if $k=1024$, one has to check $2^{1024}$ possible solutions.
Practical Data Pairs Generation.
Due to the computational challenge described above, we derive a novel data generation technique. Instead of choosing the observed vector ${\bf b}^{(j)}$ and then solving an NP-complete problem to find ${\bf x}_{{o}}^{(j)}$, it is well-known that it is much easier to find a vector $\bf b$}, given the pair $({\bf A},{\bf x})$ such that $\bf x$} is the minimizer of the QUBO problem [9]. Thus, we choose ${\bf x}_{{o}}^{(j)}$ first, and only then find a corresponding ${\bf b}^{(j)}$. This procedure allows to create data pairs while avoiding the solution of an NP-complete problem, and we consider a relaxed optimization problem, where $\bf x$} is continuous in the interval $[0,1]$:
@@
One popular way to solve this minimization problem is by the log barrier approach [36], which is obtained by adding the logarithmic terms as follows:
@@
As shown in Nocedal and Wright [36], the conditions for a minimum are:
@@
${\bf b}_{o}=-({\bf A}+{\bf A}^{\top}){\bf x}_{o}+{\frac{\mu}{{\bf x}_{o}}}-{\frac{\mu}{1-{\bf x}_{o}}},\quad{\bf b}={\bf b}_{o}+\sigma^{2}{\bf z},$ (25)
where ${\bf z}\sim N(0,{\bf I})$ is a random vector. The vector ${\bf b}_{o}$ corresponds to an approximate solution to the problem in Equation˜22, which is also the solution of the binary problem in Equation˜1. By adding a noise term $\sigma^{2}{\bf z}$ to the observed vector ${\bf b}_{o}$ in Equation˜25, the solution of the continuous problem is now different from the solution of the discrete binary problem. As explored in the previous section, the solution of the discrete problem varies little when $\sigma$ is not large. We therefore use a small number of Tabu search steps, starting from ${\bf x}_{o}$ to modify the solution if needed. While there is no guarantee that the solution is minimal, we have verified by comparison with exhaustive search solutions on small-scale problems that, indeed, we obtain the global minima every time. This approach, of choosing ${\bf x}_{o}$ rather than $\bf b$} allows us to quickly generate large amounts of data even for large-scale problems (with hundreds or thousands of decision variables), where finding the exact solution is computationally intractable. Our procedure for data generation is summarized in Algorithm 2.
@@
4:  Use up to ten Tabu search iterations to maintain efficiency, and solve Equation˜1 starting from ${\bf x}_{o}$ to obtain improved an solution $\bf x$}.
@@ 7 Numerical Experiments
Table 1: A comparison between a naive exhaustive search method, classical methods, GNNs, and our QUBO-GNN. Inference time is denoted by IT.
Size $\rightarrow$ Small ($k=10$) Medium ($k=256$) Large $(k=1024)$
Acc(%) Rel. QUBO IT(ms) Acc(%) Rel. QUBO IT(ms) Acc(%) Rel. QUBO IT(ms)
Method $\downarrow$ Objective Objective Objective
Naive.
Exhaustive 100 0 8 100 0 $10^{7}$days 100 0 $10^{302}$days
Classical.
@@
In this section, we evaluate QUBO-GNN and other baselines to understand whether our theoretical observations on the behavior of the , characterized in Section 4.2 and Section 5.1 – hold in practice.
Baselines.
@@
We consider: (i) naive methods, i.e., exhaustive search, (ii) classical solvers such as TS [18] and SAB [20], and (iii) GNN baselines, namely GCN [27], GAT[54], GIN [57], GCNII [6], and Dir-GNN [44]. We chose to compare our method with those baselines as they are representative of node classification architectures from recent years. Also, Dir-GNN is heterophily designated, thus suitable to be tested with the heterophilic behavior of QUBO solutions, as discussed in Section 5.1.
Metrics.
In our experiments, we consider three metrics:
  1. 1.